10 0/10 points I Previous Answers SerPOP5 21.P.056 My Notes Ask Your Teacher A 7.0 HF capacitor is charged by a 9.0 v battery through a resistance R. The Capacitor reaches a potential difference of 4.00 v at a time 3.00 s after charging begins. Find R. 238 Your response differs from the correct answer by more than 10%. Double check your calculations. k2 Need Help? Read
Answer #1

The potential across the capacitor after any instant of time tis, ol-e = 1-e- RC = 1 RC 3.00 R= 4 -(7x10)ln| 1- =7.298 105Ω